OpenStudy (anonymous):

Find the mean absolute deviation of the set of data 16, 20, 24, 25, 28, 29, 31 and 35.

OpenStudy (anonymous):

i thought that was how you do the mean

OpenStudy (amistre64):

the mean absolute aint the "mean"

OpenStudy (saifoo.khan):

sorry

OpenStudy (anonymous):

dnt you add all and divide by the greastest number

OpenStudy (amistre64):

sum(x-mean) ----------- N

OpenStudy (amistre64):

i get a mean of 26

OpenStudy (amistre64):

16-26 = 10 20-26 = 6 24-26 = 2 25-26 = 1 28-26 = 3 29-26 = 3 31-26 = 5 35-26 = 9 ------------ |sum| = 39 39/8 = mean absolute deviation

OpenStudy (amistre64):

4.875 :)